OpenStudy (anonymous):

how to integrate this equation? f(x)=5-(x/2)dx f(x)=int_{0}^{1}5-\frac{ x }{ 2 }dx?

OpenStudy (anonymous):

@satellite73 @hartnn @.Sam. is there any error in this question?

OpenStudy (anonymous):

fixed \[f(x)=\int\limits_{0}^{1}5-\frac{ 2 }{ x }dx?\]

OpenStudy (kc_kennylau):

You can edit the question ._.

OpenStudy (kc_kennylau):

you need to enclose the formula with a \ and a [ as well as a \ and a ]

OpenStudy (kc_kennylau):

http://screencast.com/t/vt7hEYP0

OpenStudy (kc_kennylau):

Separate it into two integrals, then use \(\large\int x^n dx=\dfrac{x^{n+1}}{n+1}\)

hartnn (hartnn):

no error in the question integral 1/x dx= ln x +c

OpenStudy (kc_kennylau):

lolz @hartnn sorry didn't notice xD

OpenStudy (kc_kennylau):

no, you have to integrate it first before plugging its values

hartnn (hartnn):

this integral will not converge as ln 0 is \(\large -\infty\)